STIMULATED EMISSION PUMPING SPECTROSCOPY OF THE X^{1}A′ STATE OF CHF

\underline{C. MUKARAKATE}, C. TAO AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; S. H. KABLE, School of Chemistry, University of Sydney, NSW 2006, Australia.

Building upon the seminal work of Suzuki and Hirota^{a}, in our first application of stimulated emission pumping spectroscopy to a halocarbene we report the spectroscopy of the X^{1}A′ state of CHF, previously examined by our group using single vibronic level emission spectroscopy.^{b}, By exciting selected rovibronic levels in the A^{1}A′′ state belonging to the progressions 2_{0}^{n}, 1_{0}^{1}2_{0}^{n}, and 2_{0}^{n}3_{0}^{1}, a variety of X^{1}A′ state levels have been observed, including several not previously observed in our emission study. The data provide precise ground state rotational constants and term energies, allowing a more complete modeling of the extensive anharmonic interactions in the X^{1}A′ state. The analysis of these interactions and the continuing search for triplet state levels will be emphasized.

aT. Suzuki and E. Hirota, J. Chem. Phys. 88, 5541 (1986).

bH. Fan, C. Mukarakate, M. Deselnicu, C .Tao, and S. A. Reid, J. Chem. Phys. 123, 014314 (2005).